# Towards Safer Chatbots: A Framework for Policy Compliance Evaluation of Custom GPTs

### v2

@@ 11 Conclusion
We presented a scalable framework for the automated evaluation of Custom GPTs’ compliance with OpenAI’s usage policies. Our results reveal that 58.7% of analyzed GPTs exhibit potential policy violations, highlighting significant gaps in the current publication and review process of the GPT store. Notably, compliance issues appear to stem primarily from the base models, such as GPT-4 and GPT-4o, rather than from user customizations. This indicates that foundational models still require improvements to better align with OpenAI’s usage policies while maintaining their utility.
Our framework, achieving an F1 score of 0.975, is an effective tool for large-scale, systematic policy compliance evaluation. Given its scalability and efficacy, we propose its integration into OpenAI’s review process for Custom GPTs prior to their publication in the GPT Store. This addition would enhance the oversight of model compliance, reducing the prevalence of unsafe or non-compliant deployments.
As future work, we plan to extend the validation of our framework to include experts across additional policy areas and sensitive domains such as healthcare, finance, and law. Although deceptive prompts were excluded from the large-scale evaluation due to annotation challenges, they remain promising for detecting complex compliance issues. We aim to refine these prompts and improve annotation consistency to incorporate them into future assessments. Additionally, we will explore advanced adversarial strategies, including multi-turn interactions and prompt injection, to uncover broader risks. These efforts will enhance the framework’s utility and contribute to safer LLM-based systems.
@@ Appendix C Operationalized OpenAI Policies
[…]
[ linewidth=1pt, roundcorner=5pt, linecolor=black, backgroundcolor=gray!10, innertopmargin=10pt, innerbottommargin=10pt, innerleftmargin=10pt, innerrightmargin=10pt ]policybox

### v3

@@ 11 Conclusion
We presented an automated and scalable method for evaluating the policy compliance of Custom GPTs deployed in the GPT Store under realistic, black-box interaction settings. Through a large-scale empirical study, we found that a substantial fraction of evaluated GPTs exhibit policy-violating behavior, revealing limitations in current publication and review practices for user-configured chatbots.
Our analysis shows that many observed violations can already be elicited from the underlying base models, such as GPT-4 and GPT-4o, using direct policy-relevant prompts. Customization typically amplifies or stabilizes these behaviors rather than introducing entirely new categories of violations, although isolated cases demonstrate that user configuration can also introduce additional risk vectors. These findings highlight the need to consider both base-model behavior and post-deployment customization when designing effective compliance and safety controls.
The validation results indicate that the proposed evaluation method can reliably reproduce human compliance judgments at the level of individual prompt–response interactions, achieving an F1 score of 0.975 under direct prompting. This demonstrates the feasibility of behavior-based, automated policy assessment at scale and suggests that such methods could complement existing manual and automated review mechanisms in large deployment ecosystems.
Looking forward, several directions remain open. Expanding validation to additional policy domains and involving annotators with expertise in regulated areas such as healthcare, finance, and law would strengthen the generality of the approach. While deceptive prompts were excluded from the large-scale evaluation due to poor annotation agreement, refining their design and improving annotation protocols may allow their inclusion in future studies. Finally, extending the method to multi-turn interactions and more adversarial prompting strategies may uncover further compliance risks that remain undetected under single-turn evaluation.
All in all, this work demonstrates that systematic, scalable policy compliance evaluation of user-configured model deployments is both necessary and technically feasible. As customizable LLM ecosystems continue to grow, such evaluation methods will be essential to support effective governance, oversight, and accountability.